OpenStudy (elizabethcastillo):

Solve the system using the elimination method. 2x+y−z=9 2x+y−z=9 −x+6y+2z=−17 −x+6y+2z=−17 5x+7y+z=4

OpenStudy (3mar):

you have three equation with three variables \[2x+y-z=9\] \[-x+6y+2z=-17\] \[5x+7y+z=4\]

OpenStudy (3mar):

to use elimination method: multiply the first and third equations by 2

OpenStudy (elizabethcastillo):

And after that step

OpenStudy (3mar):

or you can just multiply the first one by 3 and add all equations together, you will eliminate the variable z

OpenStudy (mathmale):

Eliz: It's not for us to give you the final solution; you need to get involved and find it yourself with guidance from others. What have you done so far?

OpenStudy (3mar):

for more details, you can refer to this pdf